OpenStudy (anonymous):

A stackable storage rack holds 22 cd's and costs $21. how much would it cost to buy enough racks to hold 127 dvds

OpenStudy (texaschic101):

well...do the cd's cost the same as the dvd's....and do the storage racks hold the same amount of cd's as they do dvd's ? Or is there an error in the question ?

OpenStudy (anonymous):

i ment 127 cd's not dvds @texaschic101

OpenStudy (texaschic101):

proportion time 22/21 = 127/x -- (22 cd's to $21 = 127 cd's to x dollars) cross multiply (22)(x) = (21)(127) 22x = 2667 x = 2667/22 x = 121.227 rounds to 121.23 check... 22/21 = 127/121.23 cross multiply (22)(121.23) = (21)(127) 2667.06 (rounds to 2667) = 2667 (correct) It would cost $121.23 to buy a rack that holds 127 cd's